OpenStudy (thebossofme191):

(1 point) Simplify x^13/2 · x^13/2 · x^15 · x^11 · x^26

OpenStudy (thebossofme191):

@batmano3 i need yo help plz

563blackghost (563blackghost):

Are you sure its only \(\LARGE{x^{\frac{13}{2}}}\)?

OpenStudy (thebossofme191):

yes tht is wht it showed

563blackghost (563blackghost):

Since nothing is being done to it like being multipled, divided, etc. then it would still be \(\LARGE{x^{\frac{13}{2}}}\).

OpenStudy (thebossofme191):

ohhhh *face palm*

jhonyy9 (jhonyy9):

bc you know that x^1/2 mean sqrtx so how can you rewriting this in this case using sqrt ?

jhonyy9 (jhonyy9):

x^(13/2) = x^((12+1)/2) = x^12/2 *x^1/2 yes ? so do you can continue it now ?

rebeccaxhawaii (rebeccaxhawaii):

mhm